# Comment on ”Monochromatization interaction region optics design for direct s-channel Higgs production at FCC-ee”

### v2

@@ 4 Monochromatization factor
The correct expression for $\lambda$ in the case $D_{y}^{*}=0$ can be found in [3, 5]:
@@
From Eq. (4) it follows that $\sigma_{W}$ should increase from 26.4 to 37.9 MeV for ZH4IP and from 20.2 to 41.8 MeV for ZH2IP, making the latter scheme less attractive. These values do not agree well with the simulation results presented in Table 3, which raises questions about this simulation.
A brief explanation is needed here. Simulation results can often differ from analytical formulas, since the former typically take into account nonlinear effects, beam-beam resonances, etc., which lead to perturbations of the beam distribution. However, the simulations described in [1] do not include any of these. There, the beams were represented ”as an ideal Gaussian distribution defined by the global optical parameters”. Therefore, the result should match the analytical formula, which was derived exactly for this distribution.
@@ 5 Vertical emittance
The excitation of $\varepsilon_{y}$ depends on two factors: 1) betatron coupling and 2) vertical dispersion (mainly in the arcs) which also arises due to coupling. An increase in $\varepsilon_{x}$ due to BS will lead to an increase in $\varepsilon_{y}$ due to the first factor, but not the second. To estimate this effect, it is necessary to know the ratio of the contributions of these two factors to $\varepsilon_{y0}$. However, this depends on the specific sources of X-Y coupling in the realistic magnetic lattice with errors, misalignments, etc., as well as on the methods used to correct them, and all these details are unknown for now. For a realistic assessment, it is best to use a set of simulations (with different random seeds) that take into account all the details mentioned above. If such a model is not available, then only approximate assumptions can be made.
@@ 6 Parameter optimization
The Abstract of [1] contains a phrase that is perceived as a promise that will be fulfilled: ”This measurement is significantly facilitated if the CM energy spread of $e^{+}e^{-}$ collisions can be reduced to a level comparable to the natural width of the Higgs boson, $\Gamma_{H}$ = 4.1 MeV, without substantial loss in luminosity”. And here is the result: $\sigma_{W}$ became $\sim$38 MeV, while the luminosity decreased by $\sim$2.1 times (taking into account the aforementioned factor of 1.4). Is this what we were aiming for?
As can be seen from Eq. (4), to get $\lambda\gg 1$ in a collision scheme with crossing angle and without crab crossing, we need to make $\varphi\ll 1$. To achieve this, $\sigma_{z}$ can be reduced (for example, $\sim$1.5 times) by increasing the RF voltage and $D_{x}^{*}$ should be significantly increased. The statement about a ”minimum bunch spacing of 25 ns” also raises questions. Perhaps this limitation exists at low energy (Z-pole, 45.6 GeV), where the total beam current is 3.5 times higher. However, at 62.5 GeV things become easier, so decreasing $N_{b}$ (along with increasing the number of bunches $n_{b}$) can be a useful option.
@@
The author declares that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgements
This manuscript has been authored by FermiForward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.

### v3

@@ 4 Monochromatization factor
The correct expression for $\lambda$ in the case $D_{y}^{*}=0$ and without taking into account the hourglass effect can be found in [3, 5]:
@@
From Eq. (4) it follows that $\sigma_{W}$ should increase from 26.4 to 37.9 MeV for ZH4IP and from 20.2 to 41.8 MeV for ZH2IP, making the latter scheme less attractive.
Given that the monochromatization occurs in the horizontal plane, it might seem that only the horizontal hourglass (which is negligible, since $\beta_{x}^{*}\gg\sigma_{z}$) matters. However, the vertical hourglass will also affect $\lambda$, as it leads to a change in the luminosity distribution along the azimuth, which is in some sense equivalent to a reduction in $\sigma_{z}$. Accordingly, the ”effective” Piwinski angle will decrease, and $\lambda$ will increase. To evaluate the magnitude of this effect, we used modeling with the Lifetrac beam-beam code [6]. For comparison purposes, exactly the same initial conditions were used as in [1]: BS was switched off, and the beams were represented as an ideal Gaussian distribution defined by the global optical parameters taken from Table 1 (with BS). In simulations, the vertical hourglass can be reduced to almost zero by multiple increase in $\beta_{y}^{*}$, and in this case, excellent agreement with Eq. (4) was obtained. With a nominal $\beta_{y}^{*}=1$ mm, we obtained $\sigma_{W}$ equal to 31.3 MeV for ZH4IP and 32 MeV for ZH2IP. These values do not agree well with the simulation results presented in Table 3, which raises questions about this simulation.
As can be seen, the positive effect of the vertical hourglass on $\lambda$ can be quite significant, but it weakens as $\varphi$ decreases, which is necessary for monochromatization. Besides, large hourglass would be impractical for a number of other reasons. And for $\beta_{y}^{*}\approx L_{i}$, Eq. (4) can be quite useful for estimations. It should also be noted that the precise expression for $\lambda$, taking into account the hourglass effect, is not particularly useful in our case, since there are no exact analytical formulas for calculating $\varepsilon_{x}$ considering the BS effect. The equations for $\varepsilon_{x}$ in [1] are suitable for a rough estimate, while realistic values can only be obtained through simulations, where the values of $\sigma_{W}$ and $\lambda$ can also be found directly.
@@ 5 Vertical emittance
The excitation of $\varepsilon_{y}$ depends on two factors: 1) betatron coupling and 2) vertical dispersion (mainly in the arcs) which also arises due to coupling. An increase in $\varepsilon_{x}$ due to BS will lead to an increase in $\varepsilon_{y}$ due to the first factor, but not the second. To evaluate this effect, it is necessary to know the ratio of the contributions of these two factors to $\varepsilon_{y0}$. However, this depends on the specific sources of X–Y coupling in the realistic magnetic lattice with errors, misalignments, etc., as well as on the methods used to correct them, and all these details are unknown for now. For a realistic assessment, it is best to use a set of simulations (with different random seeds) that take into account all the details mentioned above. If such a model is not available, then only approximate assumptions can be made.
@@ 6 Parameter optimization
The Abstract of [1] contains a phrase that is perceived as a promise that will be fulfilled: ”This measurement is significantly facilitated if the CM energy spread of $e^{+}e^{-}$ collisions can be reduced to a level comparable to the natural width of the Higgs boson, $\Gamma_{H}$ = 4.1 MeV, without substantial loss in luminosity”. And here is the result: $\sigma_{W}$ became 31.3 MeV, while the luminosity decreased by $\sim$2.1 times (taking into account the aforementioned factor of 1.4). Is this what we were aiming for?
As can be seen from Eq. (4), to get $\lambda\gg 1$ in a collision scheme with crossing angle and without crab crossing, we need to make $\varphi\ll 1$. To achieve this, $\sigma_{z}$ can be reduced (for example, $\sim$1.5 times) by increasing the RF voltage and $D_{x}^{*}$ should be significantly increased. The statement about a ”minimum bunch spacing of 25 ns” also raises questions. Perhaps this limitation exists at low energy (Z-pole, 45.6 GeV), where the total beam current is 3.5 times higher and the damping is 2.6 times weaker. However, at 62.5 GeV things become easier, so decreasing $N_{b}$ (along with increasing the number of bunches $n_{b}$) can be a useful option.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgements
This manuscript has been authored by Fermi Forward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. This work used resources of the Center for Research Computing and Data at Northern Illinois University.
Data availability
Data will be made available on request.